If two sets A and B have 99 elements in common ,then the number of elements common to A*B and B*A is equal to?

Grade:9

1 Answers

Arun
4 years ago
a and b having 99 elements in common . 
Then number of elements common to each of the sets a × b and b × a = 99 × 99 
= (99)² = 9801 .
Actually, you can see choose 99 elements which are common to both a and b in 99 ways separately in both the cases .
And when you will multiple these you will get the same number of elements e.g., a × b = (99)² and b × a = (99)² 
So, the answer should be 99² = 9801